Telecom Namibia is considering the purchase of a machine which cost 100 thousand dollars, and which has a lifespan of only two years, after which it has a zero scrap value. This investment, if undertaken, will generate gross return of 53 thousand dollars and 84 thousand dollars at the end of the first year and second year, respectively, after deducting all the cost except depreciation and interest rate cost. Should Telecom Namibia go ahead with this investment when the prevailing rate of interest is 18 percent? Explain


1
Expert's answer
2021-05-13T17:29:23-0400

Solution:

We use NPV to analyze the profitability of an investment.

NPV = PV of future cashflows – PV of initial investment

Under NPV calculation, depreciation is irrelevant.

First discount each future cash flow in order to get the present value of each cash flow, and then sum those present values associated with each time period.

NPV = "\sum_{t = 1}^{T}\frac{Ct}{(1+r)^{t} } - Co"

Where: Ct = Cashflow per period

             r = rate of interest

             t = time

Co = Cash outflow (initial investment)

rate of interest = 18% or 0.18

Total future cashflows for the 2 years will be as follows:

= "\frac{53,000}{(1+0.18)^{1} } + \frac{84,000}{(1+0.18)^{2} }"


= "44,915.25 + 60,327.49 = 105,242.74"

NPV = 105,242.74 – 100,000 = 5,242.74

NPV = 5,242.74


NPV is positive, which indicates that the projected earnings generated by the investment in the present exceed the expected costs, and hence will be profitable.

Therefore, Telecom Namibia should go ahead with this investment.